OpenStudy (anonymous):

Find the equation of the circle whose center and radius are given. center ( 5, 6), radius = 3

OpenStudy (anonymous):

you don't "find" it you just "write" it \[(x-h)^2+(y-k)^2=r^2\] put \(h=5,k=6, r=3\)

OpenStudy (anonymous):

let me know what you get

OpenStudy (anonymous):

wouldn't it just equal 9?

OpenStudy (anonymous):

yes, the right side would be \(3^2=9\) how about the left sides? what would that be?

OpenStudy (anonymous):

it would be 0. Because if you sub in x (which is 5) and subtract h which is 5 they would just cancel eachother out. the same for the y-k

OpenStudy (anonymous):

and 0 to the second power is 0

OpenStudy (anonymous):

ok lets go slow because i think you may be confused about what the question is asking for

OpenStudy (anonymous):

did you ever do those problems where it says "find the equation of the line given blah blah"?

OpenStudy (anonymous):

and you end up with something like \(y=2x+3\) ? does that look familiar?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so in those problems your answer was an equation that contained both an \(x\) and a \(y\)

OpenStudy (anonymous):

this is the same idea exactly, your answer will have an \(x\) and a \(y\) in it

OpenStudy (anonymous):

in other words, write the following \[(x-h)^2+(y-k)^2=r^2\] leave the \(x\) and the \(y\) in the equation only where you see a \(h\) put a \(5\) and where you see a \(k\) but a \(6\)

OpenStudy (anonymous):

*put a \(6\)

OpenStudy (anonymous):

o okay. so its (x-5)^2+(y-6)^2=3^2

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

you can write 9 on the right if you like

OpenStudy (anonymous):

now you are done, and i hope you see it was easy, but i get the feeling that it is not really clear what this means i could try to explain if you like, if not that is fine too

OpenStudy (anonymous):

no I understand! its just in a linear equation form. correct?

OpenStudy (anonymous):

well it is not a linear equation because a circle is not a line

OpenStudy (anonymous):

it is the equation of a circle not the equation for a line but suppose i told you i had an equation of a line, say \(y=2x+3\) what does that mean exactly?

OpenStudy (anonymous):

it means you have a line where y=-3/2 and im not sure how to figure out x

OpenStudy (anonymous):

* 3/2x

OpenStudy (anonymous):

ok lets back up a second math teacher often say "the equation of the line \(y=2x+3\)" but to not explain what it means, and although i don't want to slow you up it is pretty clear that you are not sure wof what it means to say that \(y=2x+3\) is the equation of a line or \((x-5)^2+(y-6)^2=9\) is the equation of a circle

OpenStudy (anonymous):

in the example of \(y=2x+3\) as the equation for a line, it means that any point on the line will fit the equation , and any point \((x,y)\) that fits the equation is on the line for example, \((7,17)\) is on the line \(y=2x+3\) because \(17=2\times 7+3\) is true

OpenStudy (anonymous):

but \((2,5)\) is not on the line \(y=2x+3\) because \(5=2\times 2+3\) is not true

OpenStudy (anonymous):

okay im getting it

OpenStudy (anonymous):

if you draw the line given by the equation \(y=2x+3\) and pick any point on the line \((a,b)\) you will see that \(b=2a+3\)

OpenStudy (anonymous):

there are an infinite number of points on the line, and all of them will fit that equation now as for the circle, it is the same idea there are an infinite number of points on the circle with center \((5,6)\) and radius \(3\)

OpenStudy (anonymous):

and any such point \((a,b)\) will make the equation \[(a-5)^2+(b-6)^2=9\] true

OpenStudy (anonymous):

okay so this equation is formatted specifically for circles? just like linear equations have their own format? I think I am getting it

OpenStudy (anonymous):

for example, \((5,3)\) is a point on the circle how do i know? because \[(5-5)^2+(3-6)^2=9\] is true

OpenStudy (anonymous):

yes, there is a specific form for the equation of a circle, which comes directly from the pythagorean theorem for a right triangle

OpenStudy (anonymous):

in your example, you had the center \((5,6)\) and the radius \(3\)

OpenStudy (anonymous):

okay! it makes a lot more sense now! I feel like I actually know what im doing now

OpenStudy (anonymous):

any point on the circle must be exactly 3 units from \((5,6)\) so by the distance formula, if \((x,y)\) is on the circle, it must fit the equation \[(x-5)^2+(y-6)^2=3^2\]

OpenStudy (anonymous):

the general form is the one i wrote at the very beginning if the center of the circle is \((h,k)\) and the radius is \(r\) then any point \((x,y)\) on the circle must satisfy the equation \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

okay! thank you so much!

OpenStudy (anonymous):

yw, hope it helped a little